OpenStudy (anonymous):

If A (with arrow above)= (6.00i-8.00j) units, B (with arrow above)= (-8.00i+ 3.00j) units, and C (with arrow above) = (26.0i+ 19.0j) units, determine a and b such that aA(with arrow above+ bB(with arrow above) + C(with arrow above) = 0. a= b=

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

\[a(6i-8j)+b(-8i+3j)=-26i-19j\] i(6a-8b)-j(8a-3b)=-26i-19j 6a-8b=-26 8a-3b=-19 a=-37/23 b=54/47